Grade 12th passAlgebra

What should be the value of an if the value of 2x2 + x – a equals 5 when x = 0?

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

From the question it is given that x = 0
We have,
2x2 + x–a = 5
a = 2x2 + x–5
Then, substitute the value of x in the equation
a = (2 × 02) + 0–5 a
= 0 + 0–5
a = -5

Thanks